OpenStudy (anonymous):

Use the data in the following table, which summarizes results from 179 pedestrian deaths that were caused by accidents. If two different deaths are randomly selected without replacement, find the probability that they both involved intoxicated driver. Pedestrian Intoxicated Yes No Driver Yes 96 40 Intoxicated No 31 12 I am not sure what I am doing wrong now. Please help me!!

OpenStudy (anonymous):

@jim_thompson5910 . something is going on. I keep doing it wrong.. Please help me. :)

jimthompson5910 (jim_thompson5910):

this is the same one from yesterday?

jimthompson5910 (jim_thompson5910):

oh but it's using different numbers

jimthompson5910 (jim_thompson5910):

P(2 intoxicated drivers) = (136/179)*(135/178)

jimthompson5910 (jim_thompson5910):

This is because there are a total of 96+40 = 136 intoxicated drivers out of 179 total

OpenStudy (anonymous):

I did it that way..

jimthompson5910 (jim_thompson5910):

and it said wrong answer?

OpenStudy (anonymous):

And the quiz answer was 0.5762 and I got 0.5772.

jimthompson5910 (jim_thompson5910):

(136/179)*(135/178) = 0.5762350134957

OpenStudy (anonymous):

It might my calculator?

jimthompson5910 (jim_thompson5910):

which rounds to 0.5762

jimthompson5910 (jim_thompson5910):

not sure how you got 0.5772

OpenStudy (anonymous):

You have 178 for the total of all of them?

jimthompson5910 (jim_thompson5910):

well it says that there are 179 of them total

jimthompson5910 (jim_thompson5910):

it drops to 178 after you pick the first driver

jimthompson5910 (jim_thompson5910):

since it says "without replacement"

OpenStudy (anonymous):

Ohh because it says two different deaths? OMG!!! that is what I messed up.. Lol!!

OpenStudy (anonymous):

Yes, I was trying to do it like that, but i didn't subst the two to have it 178..

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

Thanks again!! I really appreciate all your help..

jimthompson5910 (jim_thompson5910):

np